# Cooling and heating regions of Joule-Thomson expansion for AdS black holes: Maxwell-power-Yang-Mills and Kerr Sen black holes

### v1

@@ 5.1 Case I
Before the end of this section, if we want to talk purely about the effect of Maxwell charges as a specific result of our work, we should say that in [3], the authors found that for $q>1$ and $n=2$, the ratio $T_{i}^{min}/T_{c}$ is fully established, but in the present work, by adding the Maxwell charge to the mentioned black hole, we found that for $C\gg 1$ the Yang–Mills parameter has no effect on the ratio, and in any case, the $T_{i}^{min}/T_{c}$ is equal to $1/2$, which has the same behavior as a four-dimensional charged black hole. For $q>1$, when $C$ increases, the $T_{min}/T_{c}$ approaches the value of $1/2$. However, for $q<1$, the conditions are slightly different from the previous situation, because when $q=1/2$, $n=2$, and $Q_{1}=1$, the $T_{i}^{min}/T_{c}$ is equal to $1/2$ for all values of ”C”, which is the same as a charged black hole in four dimensions. Also, according to Table 1 to 3, it can be inferred that when $11/2<q<$ 1, with the increase of Maxwell charge, i.e. $C$, the $T_{i}^{min}/T_{c}$ only approaches the value of $1/2$ and exhibits a more similar behavior to a four-dimensional charged black hole. If for $20<q<1/$ 2, the $T_{i}^{min}/T_{c}$ becomes further or smaller than the value of $1/2$, which is contrary to a charged black hole in four dimensions. In general, when $q>1$, for the increase of Maxwell’s charge $C$, the $T_{i}^{min}/T_{c}$ is almost equal to $1/2$ and when $q=1/2$, for all values of the Maxwell’s charge, the $T_{i}^{min}/T_{c}$ is equal to $1/2$. Also, when $21<q<1/$ 2, it is close to the value of $1/2$ and finally when $01/2<q<$ 0, the values of the $T_{i}^{min}/T_{c}$ become smaller than of $1/2$.
@@ 5.2 Case II
$P(M,r_{+})=-\frac{3\left(a^{2}+r_{+}\left(2b-2\Xi^{2}M+r_{+}\right)\right)}{8\pi r_{+}(2b+r_{+})\left(a^{2}+r_{+}(2b+r_{+})\right)}.$ (26)
@@
$\begin{split}&\mathcal{X}=a^{4}(b+r_{+})+a^{2}r_{+}(4b^{2}+6br_{+}+r_{+}(2r_{+}-\Xi^{2}M))+r_{+}^{2}(2b+r_{+})(3r_{+}(b-\Xi^{2}M)+2b(b-\Xi^{2}M)+r_{+}^{2})\\ &\mathcal{Y}=2\pi r_{+}(2b+r_{+})(a^{2}+2br_{+}+r_{+})(a^{2}+r_{+}(2b+r_{+}))\\ &T(M,r_{+})=-\frac{\mathcal{X}}{\mathcal{Y}}.\end{split}$ (27)
Therefore, by using the equation (19) on can obtain,
$\begin{split}&\mathcal{A}=a^{8}(2b^{2}+2br_{+}+r_{+}^{2})+a^{6}r_{+}\big{(}16b^{3}+2b^{2}(9r_{+}+2)+br_{+}(2e^{2}M+8r_{+}+5)+r_{+}^{2}(r_{+}+2)\big{)}+a^{4}r_{+}^{2}\\ &\bigg{[}48b^{4}+4b^{3}(2\Xi^{2}M+16r_{+}+5)+4b^{2}r_{+}(8\Xi^{2}M+7r_{+}+8)+br_{+}^{2}(20\Xi^{2}M+2r_{+}+19)\\ &-r_{+}^{2}\big{(}\Xi^{2}(M-4Mr_{+})+(r_{+}-4)r_{+}\big{)}\bigg{]}+a^{2}r_{+}^{3}(2b+r_{+})\bigg{[}32b^{4}+4b^{3}(9r_{+}+4)+2b^{2}r_{+}(8\Xi^{2}M+5r_{+}+11)\\ &+br_{+}\big{(}4\Xi^{2}M(r_{+}+1)+r_{+}(11-2r_{+})\big{)}-(r_{+}-2)r_{+}^{3}\bigg{]}\\ &+(2b+1)r_{+}^{4}(2b+r_{+})^{2}\bigg{[}4b^{2}(b-\Xi^{2}M)+r_{+}^{2}(b-3\Xi^{2}M)+4br_{+}(b-\Xi^{2}M)\bigg{]}\\ &\mathcal{B}=\bigg{[}2\pi r_{+}^{2}(2b+r_{+})^{2}(a^{2}+r_{+}+2br_{+})^{2}(a^{2}+r_{+}(2b+r_{+}))^{2}\bigg{]}\\ &\mathcal{C}=3\bigg{[}a^{4}(b+r_{+})+a^{2}r_{+}\big{(}4b^{2}+6br_{+}+r_{+}(2r_{+}-\Xi^{2}M)\big{)}+r_{+}^{2}(2b+r_{+})\big{(}3r_{+}(b-\Xi^{2}M)+2b(b-\Xi^{2}M)+r_{+}^{2}\big{)}\bigg{]}\\ &\mathcal{D}=4\pi r_{+}^{2}(2b+r_{+})^{2}\big{(}a^{2}+r_{+}(2b+r_{+})\big{)}^{2}\\ &\mu=\frac{\mathcal{A}/\mathcal{B}}{\mathcal{C}/\mathcal{D}}.\end{split}$ (28)
@@
$M(P,r_{+})=\frac{\left(a^{2}+r_{+}(2b+r_{+})\right)(8\pi pr_{+}(2b+r_{+})+3)}{6\Xi^{2}r_{+}},$ (29)
$T(P,r_{+})=\frac{a^{2}\left(8\pi pr_{+}^{2}-3\right)+r_{+}^{2}(8\pi p(2b+r_{+})(2b+3r_{+})+3)}{12\pi r_{+}\left(a^{2}+2br_{+}+r_{+}\right)}.$ (30)
The $V(P)$ for the AKS black hole is calculates as follows,
$V(P)=\frac{4\pi(2b+r_{+})\left(a^{2}+r_{+}(2b+r_{+})\right)}{3\Xi^{2}}.$ (31)
So, by using the equation (19), (21), (22) and (23) one can obtain,
$\begin{split}&T_{i}=(2b+r_{+})(a^{2}+r_{+}(2b+r_{+}))\\ &\times\bigg{[}a^{4}(8\pi P_{i}r_{+}^{2}+3)+a^{2}r_{+}\bigg{(}32\pi b^{2}P_{i}r_{+}+4b(32\pi P_{i}r_{+}^{2}+3)+3(24\pi P_{i}r_{+}^{3}+r_{+}+2)\bigg{)}\\ &+16\pi(2b+1)P_{i}r_{+}^{4}(4b+3r_{+})\bigg{]}\\ &\bigg{/}12\pi r_{+}^{2}(a^{2}+2br_{+}+r_{+})^{2}\big{(}a^{2}+(2b+r_{+})(2b+3r_{+})\big{)}\end{split}$ (32)
Also, with respect to equation (30) we will have,
$\begin{split}T_{i}=\frac{a^{2}\left(8\pi r_{+}^{2}P_{i}-3\right)+r_{+}^{2}(8\pi P(2b+r_{+})(2b+3r_{+})+3)}{12\pi r_{+}\left(a^{2}+2br_{+}+r_{+}\right)}.\end{split}$ (33)
@@
We have drawn the isenthalpic curves and the inversion curve of the AdS Kerr Sen black hole for various free parameter values in each plot of Figure 4. In every subfigure, three isenthalpic curves with different M are visible, along with the corresponding inversion curve that occurs at the highest point of the isenthalpic curves. We denote each isenthalpic curve’s inversion temperature and pressure as $T_{i}$ and $P_{i}$, respectively. The isenthalpic curves are divided into two regions by the inversion curve: for $P<P_{i}$, the isenthalpic curve has a positive slope, indicating that the black hole undergoes cooling during the expansion process. With the changes in the parameters of the black hole, we found that the slope of the figures will change, and these changes are visible in each subfigure. Because an analytical solution is difficult to obtain, we used a numerical method to draw graphs. If we assume the parameter $b$ to be zero, our equations and graphs will simplify to the Kerr–AdS black holes, whose results are thoroughly discussed in [1].
Figure 4: Isenthalpic curves and inversion curve of the AdS-Ker-Sen black hole with respect to mentioned free parameters that determine in each plot.
@@
Due to the complexity of the calculations, we do not show the explicit expressions here. However, we can demonstrate the relation between the inversion temperature $T_{i}$ and the inversion pressure $P_{i}$ by using the equations given above. Figure (5a-5d) displays the inversion curves for different free parameters of the AKS black holes. The inversion curve has only one branch. The inversion temperature increases gradually with the inversion pressure, but the slope of the inversion curves becomes smaller. We can also observe some fine structures in the cases of subfigures 5. Inversion curves change with the free parameters that are specified in each plot. It was very difficult to find the analytical solution for drawing the graphs, so we used numerical solutions to draw them. Moreover, we note that due to the huge difference in the values of the horizontal and vertical graphs corresponding to the change of the free parameters, we have drawn the figures separately.
Figure 5: It shows the plot of The inversion curve for the AdS-Ker-Sen black hole with respect to free parameters that are determined in each plot with M=20
In Figure 6, we plot the Hawking temperature as a function of the horizon, which is the boundary of the black hole. The horizon can be affected by the presence of other fields or dimensions, which are represented by the free parameters in our model. We consider the free parameters constant for each plot, but we vary them across different plots to see how they influence the Hawking temperature. In each subfigure, we can observe some zero points, where the Hawking temperature becomes zero for different free parameters. This means that the black hole became extremal at these points. These zero points correspond to the divergence points of the JTC, which is a quantity that describes the temperature change when a gas expands or compresses at constant enthalpy.
Figure 6: It shows the plot of $(T-r_{+})$ for the AdS-Ker-Sen black hole with respect to free parameters that are determined in each plot
In the study of Kerr black hole[1], it was shown that the rotation parameter ”a” does not have much effect on the Joule-Thomson representation. For example, it is practically eliminated in the ratio of the $T_{i}^{min}/T_{C}$, and the value of this ratio is always a constant value of 1/2. In this paper, we found that the parameters related to AKS black hole, i.e. a and b, can play a vital role in the representation of the value of the $T_{i}^{min}/T_{C}$. The value of this ratio increases with the reduction of these parameters. For this purpose, due to the structural complexities of this black hole, we could not use the analytical method to obtain the critical points of the black hole, so we resorted to the approximate method and the numerical calculations. In the [77], the value of critical temperature and critical pressure has been calculated. We calculated some values for this black hole according to Table 4. As evident, regarding various values of free parameters, the $T_{i}^{min}/T_{C}$ is obtained, which shows the obvious effect of parameters a and b.

### v2

@@ 5.1 Case I
Before the end of this section, if we want to discuss purely the effect of Maxwell charges as a specific result of our work, we should note that in [28], the authors found that for ($q>1$ ) and ( $n=2$ ), the ratio ($T_{i}^{min}/T_{c}$) is fully established. However, in the present work, by adding the Maxwell charge to the mentioned black hole, we found that for ( $C\gg 1$ ), under the condition that we still have had black hole structures (i.e., they did not convert into a singularity system), the Yang-Mills parameter has less effect on the ratio. In any case, ($T_{i}^{min}/T_{c}$ ) is equal to ( 1/2 ), which exhibits the same behavior as a four-dimensional charged black hole. For $q>1$, when $C$ increases, the $T_{min}/T_{c}$ approaches the value of $1/2$. Also, according to Fig. (4), we can see that for q¿1 in the condition that there is no Maxwell charge, this ratio will be a constant value that is greater than 0.5, and this ratio will increase with the increase of q (in the figure as the dashed line is displayed). However, for $q<1$, the conditions are slightly different from the previous situation, because when $q=1/2$, $n=2$, and $Q=1$, the $T_{i}^{min}/T_{c}$ is equal to $1/2$ for all values of "$C"$, which is the same as a charged black hole in four dimensions. Also, according to tables (1) to (3), it can be inferred that when $11/2<q<$ 1, with the increase of Maxwell charge, i.e. $C$, the $T_{i}^{min}/T_{c}$ only approaches the value of $1/2$ and exhibits more similar behavior to a four-dimensional charged black hole. If for $20<q<1/$ 2, the $T_{i}^{min}/T_{c}$ becomes further or smaller than the value of $1/2$, which is contrary to a charged black hole in four dimensions. In general, when $q>1$, for the increase of Maxwell’s charge $C$, the $T_{i}^{min}/T_{c}$ is almost equal to $1/2$ and when $q=1/2$, for all values of the Maxwell’s charge, the $T_{i}^{min}/T_{c}$ is equal to $1/2$. Also, when $21<q<1/$ 2, it is close to the value of $1/2$ and finally when $01/2<q<$ 0, the values of the $T_{i}^{min}/T_{c}$ become smaller than of $1/2$. The proximity of the ratio $(T_{min}^{i}/T_{c})$ to 1/2 for different values of ( q ) in the AEMPYM black hole indicates a specific thermodynamic behavior related to the phase transitions and stability of the black hole. When this ratio is close to 1/2, it suggests that the black hole undergoes a Joule-Thomson expansion where the inversion temperature is approximately half of the critical temperature. This behavior can be understood as a balance between the attractive and repulsive forces within the black hole system, influenced by the Maxwell and Yang-Mills fields. The parameter ( q ) affects the strength and nature of these fields, thereby impacting the thermodynamic properties and phase transitions. In the context of the AEMPYM black hole, the Maxwell and Yang-Mills fields play a crucial role in determining the thermodynamic behavior. The parameter (q) essentially modulates the influence of these fields, leading to variations in the inversion temperature. When ( q) is greater than 1, the ratio $(T_{min}^{i}/T_{c})$ approaches 1/2 as Maxwell’s charge (C) increases, indicating a stable phase transition. For (q = 1/2), the ratio remains exactly 1/2 for all values of (C), suggesting a consistent thermodynamic behavior. However, for values of ( q ) between 0 and 1/2, the ratio decreases, moving away from 1/2, which implies a shift in the balance of forces within the black hole system, potentially leading to different phase transition characteristics.
Figure 4: The plot of $(\frac{T_{i}^{min}}{T_{c}})$ in terms of $C$. The solid blue line illustrated with $q=1.5,n=2$, and $Q_{1}=1$, solid red line denoted with $q=2,n=2$, and $Q_{1}=1$, dashed blue line represented with $q=1.5,n=2,Q_{1}=1$ and $C=0$, and dashed red line corresponds with $q=2,n=2,Q_{1}=1$ and $C=0$ for the AEMPYM black hole
@@ 5.2 Case II
$P(M,r_{+})=-\frac{3\big{(}a^{2}+r_{+}(2b-2\Xi^{2}M+r_{+})\big{)}}{8\pi r_{+}(2b+r_{+})\big{(}a^{2}+r_{+}(2b+r_{+})\big{)}}.$ (5.11)
@@
$\begin{split}&\mathcal{X}=a^{4}(b+r_{+})+a^{2}r_{+}\big{(}4b^{2}+6br_{+}+r_{+}(2r_{+}-\Xi^{2}M)\big{)}+r_{+}^{2}(2b+r_{+})\big{[}3r_{+}(b-\Xi^{2}M)+2b(b-\Xi^{2}M)+r_{+}^{2}\big{]}\\ &\mathcal{Y}=2\pi r_{+}(2b+r_{+})(a^{2}+2br_{+}+r_{+})(a^{2}+r_{+}(2b+r_{+}))\\ &T(M,r_{+})=-\frac{\mathcal{X}}{\mathcal{Y}}.\end{split}$ (5.12)
Therefore, by using Eq. (5.4) one can obtain,
$\begin{split}&\mathcal{A}=a^{8}(2b^{2}+2br_{+}+r_{+}^{2})+a^{6}r_{+}\big{(}16b^{3}+2b^{2}(9r_{+}+2)+br_{+}(2\Xi^{2}M+8r_{+}+5)+r_{+}^{2}(r_{+}+2)\big{)}+a^{4}r_{+}^{2}\\ &\bigg{[}48b^{4}+4b^{3}(2\Xi^{2}M+16r_{+}+5)+4b^{2}r_{+}(8\Xi^{2}M+7r_{+}+8)+br_{+}^{2}(20\Xi^{2}M+2r_{+}+19)\\ &-r_{+}^{2}\big{(}\Xi^{2}(M-4Mr_{+})+(r_{+}-4)r_{+}\big{)}\bigg{]}+a^{2}r_{+}^{3}(2b+r_{+})\bigg{[}32b^{4}+4b^{3}(9r_{+}+4)+2b^{2}r_{+}(8\Xi^{2}M+5r_{+}+11)\\ &+br_{+}\big{(}4\Xi^{2}M(r_{+}+1)+r_{+}(11-2r_{+})\big{)}-(r_{+}-2)r_{+}^{3}\bigg{]}\\ &+(2b+1)r_{+}^{4}(2b+r_{+})^{2}\bigg{[}4b^{2}(b-\Xi^{2}M)+r_{+}^{2}(b-3\Xi^{2}M)+4br_{+}(b-\Xi^{2}M)\bigg{]}\\ &\mathcal{B}=\bigg{[}2\pi r_{+}^{2}(2b+r_{+})^{2}(a^{2}+r_{+}+2br_{+})^{2}(a^{2}+r_{+}(2b+r_{+}))^{2}\bigg{]}\\ &\mathcal{C}=3\bigg{[}a^{4}(b+r_{+})+a^{2}r_{+}\big{(}4b^{2}+6br_{+}+r_{+}(2r_{+}-\Xi^{2}M)\big{)}+r_{+}^{2}(2b+r_{+})\big{(}3r_{+}(b-\Xi^{2}M)+2b(b-\Xi^{2}M)+r_{+}^{2}\big{)}\bigg{]}\\ &\mathcal{D}=4\pi r_{+}^{2}(2b+r_{+})^{2}\big{(}a^{2}+r_{+}(2b+r_{+})\big{)}^{2}\\ &\mu=\frac{\mathcal{A}/\mathcal{B}}{\mathcal{C}/\mathcal{D}}.\end{split}$ (5.13)
@@
$M(P,r_{+})=\frac{\left(a^{2}+r_{+}(2b+r_{+})\right)(8\pi Pr_{+}(2b+r_{+})+3)}{6\Xi^{2}r_{+}},$ (5.14)
$T(P,r_{+})=\frac{a^{2}\left(8\pi Pr_{+}^{2}-3\right)+r_{+}^{2}(8\pi P(2b+r_{+})(2b+3r_{+})+3)}{12\pi r_{+}\left(a^{2}+2br_{+}+r_{+}\right)}.$ (5.15)
The $V$ for the AKS black hole is calculated as follows[109],
$V=\frac{2\pi\big{[}a^{2}+r_{+}(2b+r_{+})\big{]}\big{[}a^{2}\big{[}r_{+}(2b+r_{+})+\ell^{2}\big{]}+2\Xi\ell^{2}r_{+}(b+r_{+})\big{]}}{3\Xi^{2}\ell^{2}r_{+}}.$ (5.16)
So, by using the Eqs. (5.4), (5.6), (5.7), and (5.8) one can obtain,
$\begin{split}&T_{i}=(2b+r_{+})(a^{2}+r_{+}(2b+r_{+}))\\ &\times\bigg{[}a^{4}(8\pi P_{i}r_{+}^{2}+3)+a^{2}r_{+}\bigg{(}32\pi b^{2}P_{i}r_{+}+4b(32\pi P_{i}r_{+}^{2}+3)+3(24\pi P_{i}r_{+}^{3}+r_{+}+2)\bigg{)}\\ &+16\pi(2b+1)P_{i}r_{+}^{4}(4b+3r_{+})\bigg{]}\\ &\bigg{/}12\pi r_{+}^{2}(a^{2}+2br_{+}+r_{+})^{2}\big{(}a^{2}+(2b+r_{+})(2b+3r_{+})\big{)}.\end{split}$ (5.17)
Also, with respect to Eq. (5.14) we will have,
$\begin{split}T_{i}=\frac{a^{2}\left(8\pi r_{+}^{2}P_{i}-3\right)+r_{+}^{2}(8\pi P_{i}(2b+r_{+})(2b+3r_{+})+3)}{12\pi r_{+}\left(a^{2}+2br_{+}+r_{+}\right)}.\end{split}$ (5.18)
@@
We have drawn the isenthalpic curves and the inversion curve of the AKS black hole for various free parameter values in each plot of Fig. (5). In every subfigure, three isenthalpic curves with different M are visible, along with the corresponding inversion curve that occurs at the highest point of the isenthalpic curves. We denote each isenthalpic curve’s inversion temperature and pressure as $T_{i}$ and $P_{i}$, respectively. The isenthalpic curves are divided into two regions by the inversion curve: for $P<P_{i}$, the isenthalpic curve has a positive slope, indicating that the black hole undergoes cooling during the expansion process. With the changes in the parameters of the black hole, we found that the slope of the figures will change, and these changes are visible in each subfigure. Because an analytical solution is difficult to obtain, we used a numerical method to draw graphs. Assuming the parameter $b$ is zero, our equations and graphs will simplify to the Kerr–AdS black holes, whose results are thoroughly discussed in [26].
Figure 5: The plot of isenthalpic and inversion curves for different values of $M$. In Fig. (5) at $a=0.5$ and $b=0.5$, in Fig. (5) at $a=0.9$ and $b=0.5$, in Fig. (5) at $a=0.9$ and $b=2.5$, and in Fig. (5) at $a=0.5$ and $b=1.5$ for the AKS black hole
@@
However, we can demonstrate the relation between the inversion temperature $T_{i}$ and the inversion pressure $P_{i}$ by using the equations given above. Figs. ((6)-(6)) displays the inversion curves for different free parameters of the AKS black holes. The inversion curve has only one branch. The inversion temperature increases gradually with the inversion pressure, but the slope of the inversion curves becomes smaller. We can also observe some fine structures in the cases of subfigures (6). Inversion curves change with the free parameters that are specified in each plot. It was very difficult to find the analytical solution for drawing the graphs, so we used numerical solutions to draw them. Moreover, we note that due to the huge difference in the values of the horizontal and vertical graphs corresponding to the change of the free parameters, we have drawn the figures separately.
Figure 6: The plot of inversion curves $(T_{i}-P_{i})$ at $a=0.5,b=0.5$ in Fig. (6), at $a=0.9,b=0.5$ in Fig. (6), at $a=0.9,b=2.5$, in Fig. (6), and at $a=0.5,b=1.5$ in Fig. (6) with respect to ($M=10,20,30$)
In Fig. (7), we plot the Hawking temperature as a function of the horizon, which is the boundary of the black hole. The horizon can be affected by the presence of other fields or dimensions, which are represented by the free parameters in our model. We consider the free parameters constant for each plot, but we vary them across different plots to see how they influence the Hawking temperature. In each subfigure, we can observe zero points, where the Hawking temperature becomes zero for different free parameters. This means that the black hole became extremal at these points. These zero points correspond to the divergence points of the JTC, which is a quantity that describes the temperature change when a gas expands or compresses at constant enthalpy. In the study of Kerr black hole[26], it was shown that the rotation parameter ”a” does not have much effect on the Joule-Thomson representation. For example, it is practically eliminated in the ratio of the $T_{i}^{min}/T_{C}$, and the value of this ratio is always a constant value of 1/2. In this paper, we found that the parameters related to AKS black hole, i.e. a and b, can play a vital role in the representation of the value of the $T_{i}^{min}/T_{C}$. The value of this ratio increases with the reduction of these parameters. For this purpose, due to the structural complexities of this black hole, we could not use the analytical method to obtain the critical points of the black hole, so we resorted to the approximate method and the numerical calculations. In the [109], the value of critical temperature and critical pressure has been calculated. We calculated some values for this black hole according to table (4). As evident, regarding various values of free parameters, the $T_{i}^{min}/T_{C}$ is obtained, which shows the obvious effect of parameters a and b. In the Kerr-AdS geometry, the rotation parameter has a minimal effect on the ratio $(T_{min}^{i}/T_{c})$ because the thermodynamic properties are primarily governed by the AdS background and the black hole’s mass. The rotation parameter, while influencing the black hole’s angular momentum and ergosphere, does not significantly alter the overall thermodynamic balance between the inversion and critical temperatures. This indicates that the rotational effects are secondary to the dominant influence of the AdS space and the black hole’s mass. The AdS background provides a stabilizing effect, ensuring that the primary thermodynamic properties remain largely unaffected by the rotation parameter. In the AKS black hole, the presence of additional fields such as the dilaton and axion fields introduces new interactions that modify the thermodynamic properties. These fields affect the black hole’s mass and angular momentum differently compared to the Kerr-AdS case. As a result, the rotation parameter has a more pronounced effect on the ratio $(T_{min}^{i}/T_{c})$. The altered behavior in the AKS case highlights the complex interplay between the black hole’s rotation, the additional fields, and the AdS background, leading to a more intricate thermodynamic structure. The dilaton and axion fields introduce additional degrees of freedom and interactions that significantly impact the black hole’s thermodynamics. These fields can alter the distribution of mass and angular momentum, leading to changes in the inversion temperature and the overall thermodynamic behavior. The rotation parameter, in this case, becomes more influential because it interacts with the dilaton and axion fields, creating a more complex and dynamic thermodynamic environment. This interplay results in a more nuanced understanding of the phase transitions and stability of the AKS black hole, as the additional fields and rotation parameters collectively shape the thermodynamic landscape.
Figure 7: The plot of temperature $T$ in terms of $r_{+}$ at $a=0.5,b=0.5$ in Fig. (7), at $a=0.9,b=0.5$ in Fig. (7), at $a=0.9,b=2.5$, in Fig. (7), and at $a=0.5,b=1.5$ in Fig. (7) for different values of $M$ for the AKS black hole